A house cat jumps from the floor and lands atop a kitchen counter that is 1.50m high. Find the minimum launch speed for the cat to successfully make such a leap.
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..2940 h=1.5m, Vi=? V = V; ² + 2gh o=v, +21-9.8)*105 V = /289-8X1.5 = 5.422 mbe